# Data-Driven Goal Recognition Design for General Behavioral Agents

### v2

@@ 4.1.1 Benchmark Domain: Grid world
In the grid world domain, agents navigate a grid with several potential goals. Take, for example, a grid world environment as illustrated in Figure 2(a). In this environment, an optimal agent starts at point ’S’ and aims for one of the goals. Spaces marked ’X’ are blocked, barring the agent’s passage. In this particular environment, the worst-case distinctiveness (wcd) is 0, as an optimal agent, following the shortest path to either goal, would reveal its intended goal on the first move due to non-overlapping paths. Our experiments are primarily focused on grid world environments with two goals for simplicity. However, our approach is applicable to scenarios with more than two goals.
Design space of environment modifications. In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces [12], also called action removal in the literature. In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
@@ 4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai. This environment is based on the popular game Overcooked. In each game, players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent. The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell. The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space. The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment. The agent can carry the movable objects, i.e. onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot). To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
Goal recognition in Overcooked-AI. The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete. For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them. In the context of goal recognition, the objective is to deduce which recipe the agent is preparing. For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
Design space of environment modifications. A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes the changing position of any object within the environment, subject to the modification being valid. This means that the change cannot result in a new design where any of the objects is unreachable by the agent. The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.
@@ 4.2 Baselines
We now describe the baseline approaches we use to compare against our approach.
@@
    Exhaustive search: This is the brute-force approach that evaluates wcd for all the environments on the search path until the minimum possible wcd is found. It is guaranteed to achieve minimum wcd. However, given the computation overhead, this approach is not applicable in most situations.
@@
    Pruned-Reduce [12]: This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
@@ 4.3 Models of Agent Behavior
    Parameterized suboptimal agent behavior. We also consider an generalized behavior model parameterized by $d(t)$. The agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
@@ 5.1 Simulations
In our simulations, we first examine how our approach compares with existing methods in standard setups found in the literature. We then showcase the generalizability of our approach by extending it to scenarios that go beyond the setups studied in the literature, including scenarios with more dynamic budget constraints, more complicated environments, and suboptimal agent behavior.
5.1.1 Standard setting
We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This scenario is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. [15]. The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by randomly allocating the blocked cells. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. We randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
@@ 5.1.1 Standard setting
(c) Shared budget
(d) Individual budget
@@
Figure 4: wcd reduction in settings with two types of modifications. We only included greedy baselines as the state-of-the-art baselines, such as Exhaustive search and Pruned-Reduce, are not applicable in these settings.
We start our experiments with a grid world of size $6\times 6$. In this simplest scenario, our approach and all baselines exhibit similar performance in wcd reduction, as depicted in Figure 3(a). For runtime comparison, our method demonstrates a significant speed advantage over exhaustive search, performing on average 10 times faster. With the maximum allowed budget of 19 modifications, our approach required only 0.2 seconds, whereas the exhaustive search took approximately 2 seconds. [^2] Detailed runtime comparisons are included in the appendix. In this simplest setting, the runtime of our approach roughly matches that of the greedy method but is 10 times faster than exhaustive search. In all other settings, our method is several orders of magnitude faster than both exhaustive search and greedy using true wcd. Although the greedy baseline becomes much faster when utilizing our predicted model for wcd, our approach remains over 3 times faster than the greedy baseline with predicted wcd for large budgets. We then expanded the grid size to $13\times 13$. In this more complex setting, both exhaustive search and the Pruned-Reduce method were unable to complete computations within an hour for an instance, leading to their exclusion from the baselines. As illustrated in Figure 3(b), our approach surpasses the greedy baselines in wcd reduction and is 3 times faster than these baselines for large budgets.
Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
Figure 6: Grid world ($6\times 6$). Incorporating suboptimal agent behavior.
5.1.2 Settings with flexible budget constraints
@@
In the literature, most works focus on a single type of environment modification (e.g., blocking a cell). Given the flexibility of our optimization framework, we extend our approach to also include ’unblocking’ blocked cells as a possible environment update. In our simulations, we examine two common cases. In the shared budget setting, the total number of blocking and unblocking actions is bounded by a given shared budget. In the individual budget setting, the number of blocking and unblocking actions is bounded by different budgets. Specifically, we allow the number of blocking actions to be 5 times the number of unblocking actions (rounded to the nearest integer). Given that there are no established baselines for this setting in the literature, we compare our results against greedy baselines. The results, as shown in Figure 4, demonstrate the effectiveness of our approach in wcd reduction. Regarding runtime, our approach is several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method with predicted wcd.
5.1.3 Complex domain and suboptimal agent behavior
@@
We consider two additional extensions to the standard setting. In the first, we evaluate the performance of our approach in a more complex problem domain: Overcooked-AI. In the second, we return to the grid world but include scenarios with suboptimal agent behavior. Both extensions utilize a grid size of $6\times 6$. For the Overcooked-AI environment, we assume optimal agent behavior and aim to explore how our approach adapts to a much richer space of environment modifications. For the suboptimal human behavior, we employ the model described in Section 4.3, utilizing a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$ and set $k=8$. For both extensions, standard approaches such as exhaustive search and Prune-Reduce are either too slow or not applicable. Therefore, we compare our results with the greedy baselines. The results for both extensions, as illustrated in Figures 6 and 6, demonstrate that our approach adapts well to both settings. Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
@@ 5.2 Real-World Human-Subject Experiments
In our simulations, we demonstrate that our approach consistently outperforms baseline methods in terms of wcd reduction and offers greater efficiency in runtime. To assess the applicability of our approach in settings where humans are the decision-makers, we conducted two sets of experiments involving human subjects. In the first experiment, our goal is to collect human behavioral data within our decision-making environments. Utilizing this data, we employ imitation learning to develop a model that accurately represents human behavior. This model is then integrated as the agent behavior model within our approach. In our second experiment, we aim to evaluate whether our approach indeed leads to environments that facilitate more effective goal recognition by human decision-makers. These experiments are approved by the Institutional Review Board (IRB) of our institution.
@@ 5.2.1 Experiment 1 : Collection of human behavioral data
In our experiment, we recruited participants to play 15 navigation games within a $6\times 6$ grid world. In each game, participants were tasked with navigating from a start position to a designated goal. At each time step, they could choose to move in one of four directions: Up, Down, Right, or Left. A game concluded when the participant reached the goal. The environments for these games were generated similarly to our simulations, with start positions, goal positions, and block positions all being randomly determined. Our objective with this setup was to leverage the collected data to develop a data-driven model of human behavior. We recruited 200 workers from Amazon Mechanical Turk for this study, paying an average hourly rate of approximately \$16 per worker.
Learning models of human behavior. The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, with approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, with around 8,800 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model for training the model. The input to the model is the current environment layout, and its output predicts the next human action. We fine-tuned the hyperparameters based on the performance on the validation dataset. We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly indicate that human behavior deviates significantly from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
Table 1: Prediction accuracy of human behavior assuming optimal behavior versus data-driven model.
Assuming Optimal Behavior Using Data-Driven Model
Training accuracy 0.7266 0.9189
Validation accuracy 0.6964 0.8136
Testing accuracy 0.7131 0.8422

### v3

@@ 4.1.1 Benchmark Domain: Grid World
In the grid world domain, agents navigate a grid with several potential goals. Consider an example in Figure 2(a): an agent starts at point ’S’ and aims for one of the goals marked ’G’. Spaces marked ’X’ are blocked. In this example, the worst-case distinctiveness (wcd) for an optimal agent is 0, as the shortest paths to the two goals do not overlap, meaning the agent’s intended goal is revealed on the very first move. Our experiments primarily focus on grid world environments with two goals for simplicity. However, our approach extends naturally to settings with more than two goals by computing the relevant GRD metric and training a corresponding predictive oracle.
Design space of environment modifications. In the context of GRD in grid world, the space of environment modification is often limited to adding blocks to spaces in the literature (Keren et al., 2014), also called action removal. In our work, we broaden the design space to also consider the unblocking of existing blocked spaces as potential modifications.
@@ 4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations in a more complex domain: Overcooked-AI. [^1] https://github.com/HumanCompatibleAI/overcooked_ai This environment is based on the popular game Overcooked, where players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the simplified setting with a single agent. The environment is represented as a grid (see Figure 2(b)), with each cell indicating the object located there. Objects may include a counter, tomato, onion, pot, dish, serving point, or empty space. The agent can only occupy empty spaces and cannot step on any other object. It can carry movable items, i.e., onions, tomatoes, and dishes, and place them on other non-space objects such as counters, serving points, or pots. To navigate the environment, the agent can move left, right, up, or down and maintains an orientation aligned with its most recent movement direction.
Goal recognition in Overcooked-AI. The agent’s goal in Overcooked-AI is defined by the recipe it needs to complete. For example, to prepare a soup with one onion and two tomatoes, the agent must collect the ingredients, place them in a pot, and cook them. In the context of goal recognition, the objective is to infer which recipe the agent is pursuing based on its behavior. For simplicity, our experiments focus on scenarios with only two possible goals or recipes.
Design space of environment modifications. A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes changing the position of any object in the environment, subject to the constraint that the modification remains valid. This means that a change cannot result in a new environment design where any of the objects is unreachable by the agent. The objective of conducting experiments in this domain is to assess whether our approach can address more complex domains.
@@ 4.2 Baselines
We compare our approach against four baselines that collectively represent the major algorithmic directions in GRD: (Optimal) exhaustive search, heuristic-enhanced search, and greedy-based approaches. For heuristic-enhanced search, we implement the Pruned-Reduce method proposed by Keren et al. (2014) as a representative example. For greedy-based approaches, we implement methods that use either the true or predicted wcd values to prioritize locally beneficial modifications.
@@
    Exhaustive search: This is the brute-force approach that evaluates wcd for all the environments on the search path until the minimum possible wcd is found. It is guaranteed to achieve minimum wcd. However, given the computational overhead, this approach is not applicable in most situations.
@@
    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
@@ 4.3 Models of Agent Behavior
    Parameterized suboptimal agent behavior. We consider a generalized behavior model parameterized by $d(t)$. The agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue & Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
@@ 5.1 Simulations
In our simulations, we first evaluate how our approach compares with existing methods in standard setups commonly found in the literature. We aim to show that our method matches or exceeds existing approaches in reducing wcd while achieving significantly better time efficiency. We then demonstrate the generalizability of our approach by applying it to scenarios that state-of-the-art methods cannot address, including those with dynamic budget constraints, more complex environments, and suboptimal agent behavior.
Standard setting. We start with settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. (2021). The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by randomly allocating the blocked cells. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. We randomly generate at least 500 environments and compare the average performance of our approach with that of baseline methods. More details of data generation and wcd predictive model training are provided in Appendix A.1 and C. During optimization for our approach, the budget cannot be directly specified, so we explored over a range of Lagrangian parameters on a log scale between 0 and 10 (detailed in Appendix subsection B.2) and record the resulting realized budgets. We use the realized budget in our experimental figures because there is a monotonic relationship: larger Lagrangian parameters correspond to stricter penalties, generally producing lower realized budgets. This allows for a clearer comparison between methods that set budgets directly and our approach, which controls budgets indirectly through Lagrangian parameters.
@@
We begin our experiments with a $6\times 6$ grid world. In this simplest setting, our approach and all baselines achieve similar performance in reducing wcd, as shown in Figure 3(a). However, our method demonstrates a significant run-time advantage over exhaustive search. In particular, our approach requires only 0.2 seconds, compared to approximately 2 seconds for exhaustive search.
We then scale up the environment to a $13\times 13$ grid, where both exhaustive search and the Pruned-Reduce method failed to complete within an hour for a single instance and were therefore excluded from the baselines. As shown in Figure 3(b), our approach outperforms the greedy baselines in reducing wcd, due to its budget-aware optimization procedure that systematically accounts for the long-term impact of each modification—making it less prone to shortsighted or myopic decisions. Additionally, our method is approximately three times faster than the greedy baselines for large budgets (with further runtime comparisons discussed later).
(a) Shared budget
(b) Individual budget
@@
Figure 4: wcd reduction in settings with two types of modifications. We only included greedy as the state-of-the-art baselines - Exhaustive search and Pruned-Reduce are not applicable in these settings.
@@
Flexible budget constraints. In the literature, most works focus on a single type of environment modification (e.g., blocking a cell). Given the flexibility of our optimization framework, we extend our approach to include ’unblocking’ blocked cells as a possible environment update. In our simulations, we examine two common cases. In “shared budget”, the total number of blocking and unblocking actions is bounded by a given shared budget. In “individual budget”, we limit the number of blocking and unblocking actions separately. Specifically, we allow the number of blocking actions to be 5 times the number of unblocking actions (rounded to the nearest integer). Given that there are no established baselines (Pruned-Reduce is limited to blocking only modifications, and Exhaustive Search does not scale) for this setting in the literature, we compare our results against greedy baselines.
The results are shown in Figure 4. Overall, they demonstrate that our approach is effective at reducing wcd even when existing state-of-the-art methods are not applicable. In addition, our gradient-based optimization outperforms the greedy approach except for small-budget cases [^2] Note that by definition of greedy approaches, the greedy approach using the true wcd is optimal when the budget is 1, since only one modification can be made. As a result, greedy methods generally perform well when the budget is very small, but our approach starts to outperform greedy with larger budgets.. This is because our method considers how to distribute the budget over time, rather than selecting the highest-impact modification at each step without considering future opportunities.
@@
Complex domain and suboptimal agent behavior. We consider two additional extensions, where existing approaches in the literature do not apply, to demonstrate the flexibility of our framework. In the first, we evaluate our approach in a more complex problem domain: Overcooked-AI. In the second, we return to the grid world but include scenarios with suboptimal agent behavior. Both extensions utilize a grid size of $6\times 6$. For the Overcooked-AI environment, we assume optimal agent behavior and aim to explore how our approach adapts to a much richer space of environment modifications. For the suboptimal human behavior, we employ the model described in Section 4.3, utilizing a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$ and set $k=8$. For both extensions, standard approaches such as exhaustive search and Prune-Reduce are either too slow or not applicable. Therefore, we compare our results with the greedy baselines.
(a) Overcooked-AI
(b) suboptimal agent.
Figure 5: Performance in other settings: (a) Overcooked-AI environment with optimal agent behavior, and (b) Grid world ($6\times 6$) with suboptimal agent behavior, demonstrating the method’s adaptability to non-optimal decision-making.
The results for both extensions, as illustrated in Figures 5(a) and 5(b), demonstrate that our approach adapts well to both settings. Our approach demonstrates a more significant advantage over baselines in Overcooked-AI than in grid world settings. The increased complexity and larger modification space in Overcooked-AI cause greedy methods to converge to short-sighted solutions, underscoring the superior scalability of our method. Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
Running time comparisons. We have demonstrated that our approach achieves greater wcd reductions compared to all baselines. Another key benefit of our method is its significantly improved runtime efficiency relative to standard approaches. This efficiency comes from (1) using a machine learning oracle to predict wcd instead of evaluating the true wcd at each time step, and (2) applying gradient-based optimization rather than the (heuristic-enhanced) exhaustive search methods used in the literature. As shown in our earlier simulations, exhaustive search methods scale poorly. In the 13x13 environment, a single instance of GRD takes more than one hour, while our approach completes in about one second. Therefore, we focus on the improvements due to the ML oracle in the following comparisons.
Table 1: Time (seconds) comparisons when using true wcd and predicted wcd. $\pm$ denotes standard error.
Setting True wcd Predicted wcd
greedy greedy our approach
[…]
Table 1 shows the time required to make modifications under a fixed budget across different settings. The results indicate our methods are generally faster and scale better in complex environments. For instance, in the complex domain, computing the true wcd takes 531 seconds for GRD, whereas our approach takes less than 1 second. Additional discussion on the time improvements is provided in Appendix A.2.
5.1.1 Ablation Study
We conducted an ablation study to better understand the specific contributions of our proposed approach to goal recognition design (GRD), particularly its ability to reduce wcd. The study isolates two key components of our method: the accuracy of the predictive model used to estimate wcd and the effectiveness of the optimization procedure in modifying the environment to improve goal recognition.
We evaluated several predictive models trained to estimate wcd and tested them in combination with different optimization strategies. These are compared against a baseline that applies a greedy search using the true wcd without relying on a predictive model. All experiments were conducted under three environment modification settings: the standard setting and two variants with flexible budget constraints (individual and shared budgets). Results for environments with grid size 13 are shown in Table 2, with additional results for grid size 6 provided in the appendix.
Impact of predictive model choice. We trained several predictive models, including Convolutional Neural Networks (CNNs), linear models, and Transformers, using the Adam optimizer (Kingma & Ba, 2015) and selected the configurations through validation. CNNs consistently achieved the lowest prediction error, and crucially, enabled more effective environment modifications when combined with either greedy or gradient-based optimization. By contrast, linear models and Transformers performed worse in predicting wcd and led to less effective downstream performance. This highlights the importance of predictive accuracy, where the quality of the predicted wcd directly influences the success of environment design.
Impact of optimization approach. We also analyzed the role of the optimization procedure given a fixed predictive model, using (i) greedy optimization using predicted wcd, and (ii) our proposed gradient-based optimization approach. Greedy optimization using predicted wcd can offer runtime benefits over using true wcd, but its effectiveness varies with prediction quality. In cases where the predictor was less accurate, the greedy method often failed to produce significant improvements in wcd. Our proposed gradient-based approach consistently improved upon the greedy method. In particular, when combined with the most accurate CNN predictor (100K samples at a learning rate of 0.001), our method achieved the lowest wcd across all three modification settings, significantly beating the greedy baseline using true wcd as well.
The best-performing configuration, CNN (100K, 0.001) combined with our gradient-based optimization, achieved the greatest reduction in wcd and outperformed all baseline approaches. This supports our core claim: accurate prediction of wcd paired with an effective optimization strategy, as adopted in our approach, is critical to designing environments that enhance goal recognition.
Table 2: Ablation results showing the impact of different predictive models and optimization methods on wcd reduction across three environment settings as described in subsection 5.1. Each row represents a predictive model trained with a specified dataset size, learning rate, and associated training/validation loss. Due to poor performance, only the best Linear and Transformer configurations are shown. For each model, we compare two optimization methods: Greedy using predicted wcd and our proposed approach, along with a baseline using greedy with true wcd. Bolded results indicate the best performance per setting. Results are for grid size 13; grid size 6 results are in Appendix B.3
Predictive Model Loss (mse) Optimization Approach wcd Reduction
Train Validation Standard Individual Shared
[…]
@@ 5.2 Real-World Human-Subject Experiments
In our simulations, we demonstrate that our approach consistently outperforms baseline methods in terms of wcd reduction and offers greater efficiency in runtime. To assess the applicability of our approach when humans are the decision-makers, we conducted two sets of human-subject experiments. In the first, we aim to collect human behavioral data in our decision-making environments. Utilizing this data, we employ imitation learning to develop a model that accurately represents human behavior. This model is then integrated as the agent behavior model within our approach. In our second experiment, we aim to evaluate whether our approach indeed leads to environments that facilitate more effective goal recognition by human decision-makers. These experiments are approved by the Institutional Review Board (IRB) of our institution.
@@ 5.2.1 Experiment 1: Collection of Human Behavioral Data
We recruited 200 Amazon Mechanical Turk workers, paying each \$1.00 for an average task time of 3.64 minutes, equivalent to an hourly rate of about \$16. Each participant is asked to play 15 navigation games in a $6\times 6$ grid world, navigating from a start position to a designated goal in each game. At each time step, they could choose to move in one of four directions: Up, Down, Right, or Left. A game concluded when the participant reached the goal. The environments for these games were generated similarly to our simulations, with start positions, goal positions, and block positions all being randomly determined. Our objective with this setup was to leverage the collected data to develop a data-driven model of human behavior.
Learning models of human behavior. The collected human data were split into training (160 workers,  70,000 user decisions), validation, and testing sets (20 workers,  8,800 decisions each). A 4-layer Multilayer Perceptron (MLP) was trained, with the environment layout as input to predict the next human action. We fine-tuned hyperparameters and chose the model architecture using validation. We compared the performance of our learned model against one that assumes optimal agent behavior i.e., taking the shortest path to the goal. Table 3 presents both models’ accuracies, showing that human behavior significantly deviates from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in GRD, particularly when humans are the decision-makers.
Table 3: Prediction accuracy of human behavior assuming optimal behavior versus using a data-driven model.
Model Train Val. Test
Assuming optimal behavior 0.7266 0.6964 0.7131
Data-driven model 0.9189 0.8136 0.8422